OpenStudy (anonymous):

Find the area of the surface generated by revolving the curve about the indicated axis. y= squareroot(5x-x^2), 0.5 <= x <= 1.5, x-axis How would I set up this integral? I know the formula but how would the integral look?

OpenStudy (tkhunny):

1) Find limits of integration. These are given x is [½.1½]. y is a little trickier, but without much difficulty, we see \(\left[\dfrac{3}{2},\dfrac{\sqrt{21}}{2}\right]\) 2) Decide Disks or Shells. It doesn't much matter in this particular problem. Pick your favorite.

OpenStudy (anonymous):

I got 5Pi is that correct?

OpenStudy (anonymous):

does anyone know if 5Pi is the answer for this?

OpenStudy (tkhunny):

Yes. That is the right answer. Did you get it by luck?

OpenStudy (anonymous):

I think so.. I set up my integral and then entered it into my calculator. So 5pi is correct? Thank you for your help

OpenStudy (tkhunny):

It's a wonderful algebra exercise. The square root is quite a mess, but the function has just what is needed to clean it up. Good work.

OpenStudy (anonymous):

Thanks again! I have another problem I am going to post up if you want to help you are more than welcome! lol